OpenStudy (anonymous):

What two numbers multiply to 192 and add to 8? Is there an easy way to figure this out?

OpenStudy (anonymous):

*might be multiplies to 192 and difference of 8 between the two numbers

OpenStudy (saifoo.khan):

xy = 192 x+y = 8

OpenStudy (anonymous):

No real solutions.

myininaya (myininaya):

x=192/y 192/y+y=8 192+y^2=8y y^2-8y+192=0 \[y=\frac{8 \pm \sqrt{64-4(192)}}{2}\] rob is right there is no such real numbers